OpenStudy (anonymous):

Let f(x) = x2(squared) − 81. Find f−1(f to the power of -1)(x)

OpenStudy (anonymous):

@Plasmataco

OpenStudy (anonymous):

A)-6 B)-3 C)3 D)6

OpenStudy (shrek_wazowski):

Like my picture :]

OpenStudy (anonymous):

I like it, been to busy trying to get work done to check chat sorry

OpenStudy (shrek_wazowski):

Ayyy lmao Tysm :3

OpenStudy (shrek_wazowski):

Also I'm not sure i'll tag a friend :0

OpenStudy (shrek_wazowski):

@KamiBug

OpenStudy (shrek_wazowski):

@Preetha

OpenStudy (shrek_wazowski):

Sorry for tagging you guys I just felt like I needed to try and help in some way :D

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle f(x)=x^2 -81 }\) and you want the inverse function \(\color{#000000}{ \displaystyle f^{-1}(x) }\) correct?

OpenStudy (solomonzelman):

I will have to dissapoint you as a mathematician, and to tell you something nice as a student doing homework. There won't be an inverse.

OpenStudy (solomonzelman):

Ok, as you probably know, the inverse of the function \(f(x)={\{\rm some~expression~with~x\}}\), is generally found by the following steps: 1) Replace f(x) with y. 2) Swipe your x and y. 3) Solve your (new) expression for y. 4) Having solved for y, write this y as \(f^{-1}(x)\) (Because this y is the inverse function).

OpenStudy (shrek_wazowski):

You're really smart xD

OpenStudy (anonymous):

i'm super confused

OpenStudy (solomonzelman):

Now, if you were to try this on a function of a form: \(\color{#000000}{ \displaystyle f(x)=x^2 -a^2 }\) Then, you would start this way: \(\color{#000000}{ \displaystyle y=x^2 -a^2 }\) (step 1) \(\color{#000000}{ \displaystyle x=y^2-a^2 }\) (step 2) \(\color{#000000}{ \displaystyle x+a^2=y^2 }\) (step 3) you can take the square root of both sides, but still don't achieve the desired result \(\color{#000000}{ \displaystyle \sqrt{x+a^2}=\sqrt{y^2} }\) and the left side is NOT just y, rather absolute value of y (via the definition of the absolute value). \(\color{#000000}{ \displaystyle \sqrt{x+a^2}=|y| }\)

OpenStudy (solomonzelman):

The most you can say is: \(\color{#000000}{ \displaystyle \left|f^{-1}(x)\right|=\sqrt{x+a^2} }\) but you can't solve for the inverse as a single function.

OpenStudy (anonymous):

I think i made this so confusing because i put the answer choices from the wrong question ':D

OpenStudy (solomonzelman):

hehe

OpenStudy (solomonzelman):

Ok, I will add the last note, to test a function for whether or not it's inverse exists. But I will add it only on a condition that you know what a one-to-one function is. Do you know?

OpenStudy (anonymous):

\[\pm \sqrt{x+81}\] \[9^{\sqrt{x}}\] \[\frac{ 1 }{ x ^{2}-81 }\] \[\frac{ x _{2} }{ 81 }\]

OpenStudy (anonymous):

no clue, it's my first week in Algebra 2

OpenStudy (solomonzelman):

Have you ever done problems finding inverse functions?

OpenStudy (solomonzelman):

Or is this the first time?

OpenStudy (anonymous):

It's my first time and i need to have it done by tonight and work in three other classes

OpenStudy (solomonzelman):

We are all busy people here. Ok, previously, I wrote the general steps of finding the iverse function. Can you please attempt an example problem: \(f(x)=x+4\)?

OpenStudy (solomonzelman):

The inverse of the function \(f(x)=\rm \{some~expression~with~x\}\), is generally found by the following steps: 1) Replace f(x) with y. 2) Swipe your x and y. 3) Solve your (new) expression for y. 4) Having solved for y, write this y as f\(^{-1}\)(x). (since this y is inverse function).

OpenStudy (anonymous):

But don't i have to know what X is ?

OpenStudy (solomonzelman):

Please try to preform these steps on f(x)=x+4. No, you don't have to know what x is. It could be anything, since it is the independent variable.

OpenStudy (anonymous):

So, is sit y=x-4 ?

OpenStudy (anonymous):

is it*

OpenStudy (solomonzelman):

that would be it. and then re-write this as \(f^{-1}(x)=x-4\). So, \(f^{-1}(x)\) instead of the y, since it is the inverse function.

OpenStudy (solomonzelman):

Ok, can you do one more for me please? f(x)=5x-2

OpenStudy (anonymous):

Would you change that to y=5x-2 and then do 5x=y-2 or x=5y-2 ??

OpenStudy (anonymous):

????

OpenStudy (solomonzelman):

f(x)=5x-2 y=5x-2 x=5y-2 x+2=5y (x+2)/5=y So, \(f^{-1}(x)=(x+2)/5\)

OpenStudy (solomonzelman):

Ok ?

OpenStudy (anonymous):

uh, ohkai

OpenStudy (solomonzelman):

Ok, now let's try f(x)=x^2-81

OpenStudy (solomonzelman):

Can you do the first 2 steps for me please?

OpenStudy (anonymous):

I think i got the answer, but i don't understand the \[\pm\] symbol

OpenStudy (solomonzelman):

Ok, can you show your work?

OpenStudy (solomonzelman):

I will tell you where you are having a problem

OpenStudy (anonymous):

\[y=\sqrt{81+x}\] is what i got

OpenStudy (solomonzelman):

That is not correct.

OpenStudy (anonymous):

oh, sorry yeah gimme one second to get it all in

OpenStudy (anonymous):

why not ?

OpenStudy (solomonzelman):

sure, take your time.

OpenStudy (solomonzelman):

Can you type your work. All of it, please.

OpenStudy (solomonzelman):

I want to see all steps

OpenStudy (anonymous):

\[y=x ^{2}-81\] \[x=y ^{2}-81\] +81 +81 \[x+81=y ^{2}\] \[\sqrt{x+81}=\sqrt{y ^{2}}\] \[y=\sqrt{x+81}\]

OpenStudy (solomonzelman):

I am sure you are familiar with the absolute value function, (right?) So, there is a defintion for the absolute value: \(\color{#000000}{ \displaystyle \sqrt{z^2}=\left|z\right| }\)

OpenStudy (solomonzelman):

So, actually, after you take the square root of both sides to get: \(\color{#000000}{ \displaystyle \sqrt{x+81}=\sqrt{y^2} }\) And by the definition of the absolute value you are going to get: \(\color{#000000}{ \displaystyle \sqrt{x+81}=|y| }\) You know that in general, if \(\color{#000000}{ \displaystyle |z|=5 }\), then \(z=\pm 5\). By this convention, you have, \(\color{#000000}{ \displaystyle \pm \sqrt{x+81}=y }\) And since y is the inverse function you write, \(\color{#000000}{ \displaystyle f^{-1}(x)= \pm \sqrt{x+81} }\)

OpenStudy (solomonzelman):

There is a technical error with saying that this is a valid inverse and with saying that the inverse of x^2-81 exists, because both of these statements are wrong!

OpenStudy (anonymous):

well, i'm gonna go with that one because none of the above isn't an option and i still have too much other work to do. But i understand a lot better, thank you (:

OpenStudy (solomonzelman):

Yes, go with that one. I was just going to say that don't worry about this technicality now, except for that I didn't have time because I lagged the 51st time.

OpenStudy (solomonzelman):

Enjoy, and, best of luck to your mathematics!

OpenStudy (anonymous):

thank you (:

OpenStudy (mathmale):

It's true that the given function doesn't have an inverse (if you were to graph y=x^2 - 81, you 'd see a parabola that opens upward. Drawing a horiz. line thru this graph would intersect the graph in 2 places, which means that y = x^2 - 81 has no inverse. However, you could restrict the domain of y = x^2 - 81 to [0, infinity), in which case this derived function WOULD have an inverse.

OpenStudy (anonymous):

0.o my brain hurts. I'm gonna post another question because i'm stupid xD

OpenStudy (mathmale):

Sorry about your brain hurting. But kindly stop using "I'm stupid" as a reason or an excuse for anything. Why shoot yourself in the foot if you don't need or want to ? Ouch.

OpenStudy (damonte212):

..

OpenStudy (mathmale):

Please repost your latest question, but separately from this post. Thx.